12. For an F distribution, find the following: (a) fo.25,5,10 (b) fo.10.c)o.05,8,15 (d) o.75,5,10 (e)fo.so, (f)fo.95,8,15 0.10,24,9 0.90.24,9
Answer #1

(a) 25, 5, 10 1.5853 ; From Excel "=FINV(0.25,5,10)"

(b) fo·25.5.10 = 2.2768 ; From Excel "=FINV(0.10,24,9)"

(c) 10.25, 5.10 2.6408 ; From Excel "=FINV(0.05,8,15)"

(d) 25, 5, 10 = 0.5291 ; From Excel "=FINV(0.75,5,10)"

(e) ん25,5,10 0.5246 ; From Excel "=FINV(0.90,24,9)"

(a) f0. 25.5.10 = 0.3107 ; From Excel "=FINV(0.95,8,15"